Please explain how to do this so I can complete the other questions - Thank you!Express your answer as an integer. PO ALQ O a ? 3.42 pm Submit Request Answer
Answer #1

we know that, 1 um= 104 cm x= 3.42 4m = 3.42×10 cm 3042X10-4cm. . : 2923.97 em [3.42 um=2923.97cm Angwe know that, 1 um= 104 cm x= 3.42 4m = 3.42×10 cm 3042X10-4cm. . : 2923.97 em [3.42 um=2923.97cm Ang